Force homework? I'm really having a problem with these questions on my homework, and could really use some help. Explanations would be really appreciated as well.

1. When you push a 1.80 kg book resting on a tabletop, it takes 1.70 N to start the book sliding. What is the frictional force exerted on the book when you push on it with a force of 0.65 N?

2. To move a large crate across a rough floor, you push on it with a force F at an angle of 21 degrees below the horizontal. Find the acceleration of the crate if the applied force is 380 N, the mass of the crate is 32 kg, and the coefficient of kinetic friction is 0.38.
Answer #1

part 1 )

it will be the same as applied force because book will not be slide by 0.65 N

friction force = 0.65 N

part 2 )

Fnet = Fcos(21) - u(mg-Fsin(21)) = 0

Fcos(21) + uFsin(21) = umg

F[cos(21) + usin(21)] = umg

F = umg / [cos(21) + usin(21)]

F = [.38 x 32kg x 9.8m/s^2] / (1.07)

F = 111.37N

F = ma

111.37 / 32 = 3.48 m/s^2